OpenStudy (anonymous):

Gravitational force is affected by:

OpenStudy (anonymous):

the object's mass, and the distance between two objects

OpenStudy (anonymous):

I think Mass and Earth.

OpenStudy (king.void.):

Options?

OpenStudy (abhisar):

Gravitational force is given by the formula \(\sf \boxed{\huge F_g=\frac{G\times M_1\times M_2}{R^2}}\) In the above formula, G is a constant called as Universal gravitation Constant. \(\sf M_1~and~M_2\) are masses of the two objects and R is the distance between them. Since, the masses and the distance between them are the only two variables, the gravitational force should depend upon?